# Neural Network-based Universal Formulas for Control

### v1

@@ 3 A Universal Formula for an Arbitrary Number of Affine Constraints
$\nabla^{2}J_{i,{\mathbf{p}}}({\mathbf{k}})$ $=\frac{-1}{(A_{i}\!+\!{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{3}}\big{(}(A_{i}\!+\!{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\mathbf{I}_{n}\!-\!({\mathbf{k}}{\mathbf{B}}_{i}^{\top}\!+\!{\mathbf{B}}_{i}{\mathbf{k}}^{\top})(A_{i}\!+\!{\mathbf{B}}_{i}^{\top}{\mathbf{k}})+(\lVert{\mathbf{B}}_{i}\rVert^{2}\!+\!\lVert{\mathbf{k}}\rVert^{2}){\mathbf{B}}_{i}{\mathbf{B}}_{i}^{\top}\big{)}.$
Let us then show that the matrix $\Gamma=(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\mathbf{I}_{n}-({\mathbf{k}}{\mathbf{B}}_{i}^{\top}+{\mathbf{B}}_{i}{\mathbf{k}}^{\top})(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})+(\lVert{\mathbf{B}}_{i}\rVert^{2}+\lVert{\mathbf{k}}\rVert^{2}){\mathbf{B}}_{i}{\mathbf{B}}_{i}^{\top}$ is positive definite, i.e., ${\mathbf{x}}^{\top}\Gamma{\mathbf{x}}>0$ whenever ${\mathbf{x}}\in\mathbb{R}^{m}\backslash\{\mathbf{0}_{m}\}$. This is true iff
$(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\lVert{\mathbf{x}}\rVert^{2}-2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{k}}^{\top}{\mathbf{x}})+(\lVert{\mathbf{B}}\rVert_{i}^{2}+\lVert{\mathbf{k}}\rVert^{2})({\mathbf{B}}_{i}^{\top}{\mathbf{x}})^{2}>0,$
for all ${\mathbf{x}}\in\mathbb{R}^{m}\backslash\{\mathbf{0}_{m}\}$. Let $z={\mathbf{B}}_{i}^{\top}{\mathbf{x}}$. We want to show that the scalar quadratic function in $z$ given by $(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\lVert{\mathbf{x}}\rVert^{2}-2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{k}}^{\top}{\mathbf{x}})z+(1+\lVert{\mathbf{k}}\rVert^{2})z^{2}$ is strictly positive for all ${\mathbf{x}}\in\mathbb{R}^{m}\backslash\{\mathbf{0}_{m}\}$. Note that the discriminant of the quadratic is
@@
By the Cauchy-Schwartz inequality, $({\mathbf{k}}^{\top}{\mathbf{x}})^{2}\leq\lVert{\mathbf{k}}\rVert^{2}\lVert{\mathbf{x}}\rVert^{2}$. Therefore, $({\mathbf{k}}^{\top}{\mathbf{x}})^{2}-\lVert{\mathbf{k}}\rVert^{2}\lVert{\mathbf{x}}\rVert^{2}-\lVert{\mathbf{k}}\rVert^{2}\lVert{\mathbf{B}}_{i}\rVert^{2}<0$ unless ${\mathbf{B}}_{i}=\mathbf{0}_{m}$ (in which case $(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\lVert{\mathbf{x}}\rVert^{2}-2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{k}}^{\top}{\mathbf{x}})z+(1+\lVert{\mathbf{k}}\rVert^{2})z^{2}>0$ for ${\mathbf{x}}\neq\mathbf{0}_{n}$) or ${\mathbf{x}}=\mathbf{0}_{m}$. This implies that $\Delta<0$ unless ${\mathbf{x}}=\mathbf{0}_{m}$ and therefore the quadratic $(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\lVert{\mathbf{x}}\rVert^{2}-2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{k}}^{\top}{\mathbf{x}})z+(1+\lVert{\mathbf{k}}\rVert^{2})z^{2}$ is strictly positive for all ${\mathbf{x}}\in\mathbb{R}^{m}\backslash\{\mathbf{0}_{m}\}$. Hence, $\Gamma$ is positive definite, which means that $\nabla^{2}J_{i,{\mathbf{p}}}({\mathbf{k}})$ is positive definite for all ${\mathbf{k}}\in\mathcal{K}$, and $J_{i,{\mathbf{p}}}$ is strictly convex in $\mathcal{K}$. Since $J_{{\mathbf{p}}}({\mathbf{k}})=\sum_{i=1}^{N}J_{i,{\mathbf{p}}}({\mathbf{k}})$, $J_{{\mathbf{p}}}$ is the sum of strictly convex functions and is therefore also strictly convex in $\mathcal{K}_{{\mathbf{p}}}$. ∎
Since $J_{{\mathbf{p}}}$ is strictly convex in $\mathcal{K}_{{\mathbf{p}}}$ by Proposition 3.1, it has a unique minimizer in $\mathcal{K}_{{\mathbf{p}}}$. Let
@@
Since ${\mathbf{p}}\in\mathcal{P}$, $\mathcal{K}_{{\mathbf{p}}}\neq\emptyset$. First note that since $\lim\limits_{\lVert{\mathbf{k}}\rVert\to\infty}J_{{\mathbf{p}}}({\mathbf{k}})=\infty$, $k^{*}({\mathbf{p}})$ is finite for all ${\mathbf{p}}$. Since $J_{{\mathbf{p}}}$ is strictly convex, continuous in $\mathcal{K}_{{\mathbf{p}}}$ and $k^{*}({\mathbf{p}})$ is its unique minimizer, there exist neighborhoods $\mathcal{N}_{1}$, $\mathcal{N}_{2}$ of $k^{*}({\mathbf{p}})$ and a constant $M>0$ such that if ${\mathbf{k}}\in\mathcal{N}_{1}$ then $J_{{\mathbf{p}}}({\mathbf{k}})<M$, and if ${\mathbf{k}}\notin\mathcal{N}_{2}$ then $J_{{\mathbf{p}}}({\mathbf{k}})>2M$. Now, since $J_{{\mathbf{p}}}$ is continuous with respect to ${\mathbf{p}}$, there exists a neighborhood $\mathcal{N}_{{\mathbf{p}}}$ of ${\mathbf{p}}$ such that for any $\bar{{\mathbf{p}}}\in\mathcal{N}_{{\mathbf{p}}}$, we have that if ${\mathbf{k}}\in\mathcal{N}_{1}$ then $J_{\bar{{\mathbf{p}}}}({\mathbf{k}})<\frac{3M}{2}$, and if ${\mathbf{k}}\notin\mathcal{N}_{2}$ then $J_{{\mathbf{p}}}({\mathbf{k}})>\frac{3M}{2}$. Hence, $k^{*}(\bar{{\mathbf{p}}})\in\mathcal{N}_{1}$ for all $\bar{{\mathbf{p}}}\in\mathcal{N}_{{\mathbf{p}}}$. Now, define the function $J:\mathcal{N}_{1}\times\mathcal{N}_{1}$ as $J({\mathbf{p}},{\mathbf{k}})=J_{{\mathbf{p}}}({\mathbf{k}})$. Note that since $k^{*}$ is the minimizer of $J_{{\mathbf{p}}}$, it satisfies the equation $\frac{\partial J}{\partial{\mathbf{k}}}(\bar{{\mathbf{p}}},k^{*}(\bar{{\mathbf{p}}}))=\mathbf{0}_{m}$ for all $\bar{{\mathbf{p}}}\in\mathcal{N}_{1}$. Additionally, note that since $J_{{\mathbf{p}}}$ is strictly convex as shown in Proposition 3.1, $\frac{\partial^{2}J}{\partial^{2}{\mathbf{k}}}(\bar{{\mathbf{p}}},k^{*}(\bar{{\mathbf{p}}}))$ is non-singular. Now, by the Implicit Function Theorem (Dontchev and Rockafellar, 2014, Proposition 1B.5), $k^{*}$ is $\mathcal{C}^{\infty}$ at ${\mathbf{p}}$. This argument is valid for any ${\mathbf{p}}\in\mathcal{P}$. ∎
@@
(Connection with Sontag’s universal formula): A noteworthy property of the controller $u^{*}$ is that when $N=1$, $m=1$, and the constraint corresponds to a CLF constraint, i.e., $a_{1}({\mathbf{x}})=\nabla V({\mathbf{x}})^{\top}f({\mathbf{x}})+W({\mathbf{x}})$, $b_{1}({\mathbf{x}})=\nabla V({\mathbf{x}})^{\top}g({\mathbf{x}})$ for some CLF $V$ and positive definite function $W$, then $k^{*}$ equals the celebrated Sontag’s universal formula (Sontag, 1989) for stabilization. Indeed, the minimizer of $J_{{\mathbf{p}}}$ can be found by solving the nonlinear equation $\nabla J_{{\mathbf{p}}}(k)=-\frac{k}{A_{1}+B_{1}k}+\frac{B_{1}^{2}+k^{2}}{2(A_{1}+B_{1}k)^{2}}B_{1}=0$. By solving the resulting quadratic equation, one obtains Sontag’s universal formula:
$k^{*}(A_{1},B_{1})=\begin{cases}-\frac{A_{1}+\sqrt{A_{1}^{2}+B_{1}^{4}}}{B_{1}},\ &\text{if}\ B_{1}\neq 0,\\ 0,&\text{otherwise}.\qquad\qquad\qquad\qquad\penalty 9999\hbox{}\nobreak\hfill\quad\hbox{$\bullet$}\end{cases}$
@@
where $J({\mathbf{x}},{\mathbf{u}})=J_{{\mathbf{x}}}({\mathbf{u}})$ and $\tau>0$ is a design parameter. Then, by using singular perturbation theory (Khalil, 2002, Chapter 11), one can show that for sufficiently large $\tau$, the evolution of the state variable ${\mathbf{x}}$ according to (6) can be made arbitrarily close to the evolution of the state variable for $\dot{{\mathbf{x}}}=f({\mathbf{x}})+g({\mathbf{x}})u^{*}({\mathbf{x}})$. $\bullet$
@@
$J_{{\mathbf{p}}}^{{\mathbf{w}}}({\mathbf{k}})=-\sum_{i=1}w_{i}\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}+\lVert{\mathbf{k}}\rVert^{2}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}.$
@@ 4 Neural Network Approximation of the Universal Formula
$\tilde{J}_{{\mathbf{q}}({\mathbf{p}})}({\mathbf{k}})=-M\sum_{i=1}^{N}\frac{\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}}{M^{2}}+\frac{\lVert{\mathbf{k}}\rVert^{2}}{M^{2}}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}=-\frac{1}{M}\sum_{i=1}^{N}\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}+\lVert{\mathbf{k}}\rVert^{2}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}=J_{{\mathbf{p}}}({\mathbf{k}}).$
@@
Lemma 4.1 shows that by approximating $\tilde{k}^{*}$ with values of ${\mathbf{q}}$ in $\mathcal{T}=\left([-1,1]^{N}\times\mathcal{B}_{1}(\mathbf{0}_{m})^{N}\right)\cap\mathcal{P}\times[0,1]$, we can recover any value of $k^{*}$. Therefore, our approach consists in approximating the minimizer of $\tilde{J}_{{\mathbf{q}}}$ with a NN for ${\mathbf{q}}$ in the set $\mathcal{T}$. Since $\mathcal{P}$ is an open set, $\mathcal{T}$ is not compact. Given that the universal approximation theorem of neural networks (Cybenko, 1989) is only valid in compact sets, it does not apply to $\mathcal{T}$. However, for any compact subset of $\mathcal{T}$ (which we can take to be arbitrarily close to $\mathcal{T}$), the universal approximation theorem ensures that a NN with a sufficiently large number of parameters can approximate $k^{*}$ arbitrarily well. Note also that by choosing the activation functions of the NN to be smooth, the resulting approximation of $k^{*}$ is $\mathcal{C}^{\infty}(\mathcal{P})$, and therefore the approximation of the controller $u^{*}$ is $\mathcal{C}^{l}(\mathcal{P})$.
@@
(Applicability for multiple controllers): Note that once we have obtained an approximation of $k^{*}$, we can use it in many different instantiations of the functions $a_{1},\ldots,a_{N},b_{1},\ldots,b_{N}$, i.e., in a variety of different control problems. The NN only depends on the number of constraints $N$ and the dimension of the input $m$, but remarkably, does not depend on $n$, the dimension of the state. In fact, given a NN trained with a given input dimension $m$ and number of constraints $N$, all control problems with input dimension at most $m$ and at most $N$ constraints can be solved with the same NN (in case the input dimension is strictly less than $m$ or the number of constraints is strictly less than $N$ one can simply assign the coefficients of the additional inputs or constraints so that they are trivially satisfied). In particular, this implies the remarkable fact that all safe stabilization problems for systems with a given input dimension can be solved with the same NN approximation. $\bullet$
@@
(Robustness): If one of the inequalities corresponds to a CLF inequality (2), and the corresponding CLF is an input-to-state stable (ISS) Lyapunov function for the system (1) (cf. (Sontag, 2008, Section 3.3)), then if we let $\hat{k}$ be an approximated version of the controller $k^{*}$ (like the one obtained by approximating it with a NN), and the bound $\big{\lVert}\hat{k}({\mathbf{x}})-k^{*}({\mathbf{x}})\big{\rVert}\leq\sigma$ holds for all ${\mathbf{x}}\in\mathcal{X}$, with $\mathcal{X}$ some compact set, we have
$\nabla V({\mathbf{x}})^{\top}(f({\mathbf{x}})+g({\mathbf{x}})\hat{k}({\mathbf{x}}))=\nabla V({\mathbf{x}})^{\top}(f({\mathbf{x}})+g({\mathbf{x}})k^{*}({\mathbf{x}}))+\gamma(\sigma),$
for some class $\mathcal{K}_{\infty}$ function $\gamma$. A similar result holds if the inequality is a CBF inequality (3) and the CBF defining it is an input-to-state safe (ISSf) barrier function for the system (1(Kolathaya and Ames, 2018, Theorem 1). Therefore, CLF and CBF inequalities are robust to approximation errors induced by the NN and therefore the NN-based controller, when applied to safety and stability tasks, makes the closed-loop system close to safe and stable, respectively. $\bullet$
@@
(Hard-constrained NN): We should also point out that the recently introduced HardNet method (Min and Azizan, 2025) shows that, by including a projection step at the output layer of the NN, the NN predictions can be guaranteed to satisfy a set of affine constraints. Although the universal approximation guarantees are retained, the projection operation in the last layer makes the controller obtained from it locally Lipschitz but not $\mathcal{C}^{l}(\mathcal{F})$, for $l\geq 1$. $\bullet$

### v2

@@ 3 A Universal Formula for an Arbitrary Number of Affine Constraints
$\nabla^{2}J_{i,{\mathbf{p}}}({\mathbf{k}})=\frac{-\Gamma}{(A_{i}\!+\!{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{3}},$
[…]
$(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})^{2}\lVert{\mathbf{x}}\rVert^{2}-2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{B}}_{i}^{\top}{\mathbf{k}})({\mathbf{k}}^{\top}{\mathbf{x}})+$
$(\lVert{\mathbf{B}}\rVert_{i}^{2}+\lVert{\mathbf{k}}\rVert^{2})({\mathbf{B}}_{i}^{\top}{\mathbf{x}})^{2}>0,$
[…]
@@
[…]
Since $J_{{\mathbf{p}}}$ is strictly convex in $\mathcal{K}_{{\mathbf{p}}}$ by Proposition 3.1, and $J_{{\mathbf{p}}}$ goes to infinity as it approaches the boundary of $\mathcal{K}_{{\mathbf{p}}}$, it follows that $J_{{\mathbf{p}}}$ has a unique minimizer in $\mathcal{K}_{{\mathbf{p}}}$. Let
@@
Since ${\mathbf{p}}\in\mathcal{P}$, $\mathcal{K}_{{\mathbf{p}}}\neq\emptyset$. First note that since $\lim\limits_{\lVert{\mathbf{k}}\rVert\to\infty}J_{{\mathbf{p}}}({\mathbf{k}})=\infty$, $k^{*}({\mathbf{p}})$ is finite for all $\mathbf{p}$}. Since $J_{{\mathbf{p}}}$ is strictly convex, continuous in $\mathcal{K}_{{\mathbf{p}}}$ and $k^{*}({\mathbf{p}})$ is its unique minimizer, there exist neighborhoods $\mathcal{N}_{1}$, $\mathcal{N}_{2}$ of $k^{*}({\mathbf{p}})$ and a constant $M>0$ such that if ${\mathbf{k}}\in\mathcal{N}_{1}$ then $J_{{\mathbf{p}}}({\mathbf{k}})<M$, and if ${\mathbf{k}}\notin\mathcal{N}_{2}$ then $J_{{\mathbf{p}}}({\mathbf{k}})>2M$. Now, since $J_{{\mathbf{p}}}$ is continuous with respect to $\mathbf{p}$}, there exists a neighborhood $\mathcal{N}_{{\mathbf{p}}}$ of $\mathbf{p}$} such that for any $\bar{{\mathbf{p}}}\in\mathcal{N}_{{\mathbf{p}}}$, we have that if ${\mathbf{k}}\in\mathcal{N}_{1}$ then $J_{\bar{{\mathbf{p}}}}({\mathbf{k}})<\frac{3M}{2}$, and if ${\mathbf{k}}\notin\mathcal{N}_{2}$ then $J_{{\mathbf{p}}}({\mathbf{k}})>\frac{3M}{2}$. Hence, $k^{*}(\bar{{\mathbf{p}}})\in\mathcal{N}_{1}$ for all $\bar{{\mathbf{p}}}\in\mathcal{N}_{{\mathbf{p}}}$. Now, define the function $J:\mathcal{N}_{1}\times\mathcal{N}_{1}\to\mathbb{R}$ as $J({\mathbf{p}},{\mathbf{k}})=J_{{\mathbf{p}}}({\mathbf{k}})$. Note that since $k^{*}$ is the minimizer of $J_{{\mathbf{p}}}$, it satisfies the equation $\frac{\partial J}{\partial{\mathbf{k}}}(\bar{{\mathbf{p}}},k^{*}(\bar{{\mathbf{p}}}))=\mathbf{0}_{m}$ for all $\bar{{\mathbf{p}}}\in\mathcal{N}_{1}$. Additionally, note that since $J_{{\mathbf{p}}}$ is strictly convex as shown in Proposition 3.1, $\frac{\partial^{2}J}{\partial^{2}{\mathbf{k}}}(\bar{{\mathbf{p}}},k^{*}(\bar{{\mathbf{p}}}))$ is non-singular. Now, by the Implicit Function Theorem [31, Proposition 1B.5], since $J$ is $\mathcal{C}^{\infty}$, $k^{*}$ is also $\mathcal{C}^{\infty}$ at $\mathbf{p}$}. This argument is valid for any ${\mathbf{p}}\in\mathcal{P}$. ∎
@@
(Connection with CLF universal formula): A noteworthy property of the controller $u^{*}$ is that when $N=1$, $m=1$, and the constraint corresponds to a CLF constraint, i.e., $a_{1}({\mathbf{x}})=\nabla V({\mathbf{x}})^{\top}f({\mathbf{x}})+W({\mathbf{x}})$, $b_{1}({\mathbf{x}})=\nabla V({\mathbf{x}})^{\top}g({\mathbf{x}})$ for some CLF $V$ and positive definite function $W$, then $k^{*}$ equals the well-known universal formula [4] for stabilization. Indeed, the minimizer of $J_{{\mathbf{p}}}$ can be found by solving the nonlinear equation $\nabla J_{{\mathbf{p}}}(k)=-\frac{k}{A_{1}+B_{1}k}+\frac{B_{1}^{2}+k^{2}}{2(A_{1}+B_{1}k)^{2}}B_{1}=0$. By solving the resulting quadratic equation, one obtains the CLF universal formula:
$k^{*}(A_{1},B_{1})=\begin{cases}-\frac{A_{1}+\sqrt{A_{1}^{2}+B_{1}^{4}}}{B_{1}},\ &\text{if}\ B_{1}\neq 0,\\ 0,&\text{otherwise}.\qquad\penalty 9999\hbox{}\penalty 10000\hfill\quad\hbox{$\bullet$}\end{cases}$
@@
where $J({\mathbf{x}},{\mathbf{u}})=J_{{\mathbf{x}}}({\mathbf{u}})$ and $\tau>0$ is a design parameter. Then, by using singular perturbation theory [32, Chapter 11], one can show that for sufficiently large $\tau$, the evolution of the state variable $\mathbf{x}$} according to (6) can be made arbitrarily close to the evolution of the state variable for $\dot{{\mathbf{x}}}=f({\mathbf{x}})+g({\mathbf{x}})u^{*}({\mathbf{x}})$.  $\bullet$
@@
$J_{{\mathbf{p}}}^{{\mathbf{w}}}({\mathbf{k}})=-\sum_{i=1}^{N}w_{i}\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}+\lVert{\mathbf{k}}\rVert^{2}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}.$
@@
The following result shows that for safe stabilization problems, even though $u^{*}$ is not defined at the origin, trajectories in its neighborhood converge to it.
Proposition 3.6.
(Convergence to the origin): Let $V:\mathbb{R}^{n}\to\mathbb{R}$ be a CLF, $h:\mathbb{R}^{n}\to\mathbb{R}$ be a CBF of a safe set $\mathcal{C}\subset\mathbb{R}^{n}$. Let $N=2$ and define $a_{1}({\mathbf{x}})=\nabla V({\mathbf{x}})^{\top}f({\mathbf{x}})+W({\mathbf{x}})$, $b_{1}({\mathbf{x}})=g({\mathbf{x}})^{\top}\nabla V({\mathbf{x}})$, $a_{2}({\mathbf{x}})=-\nabla h({\mathbf{x}})^{\top}f({\mathbf{x}})-\alpha h({\mathbf{x}})$, $b_{2}({\mathbf{x}})=-g({\mathbf{x}})^{\top}\nabla h({\mathbf{x}})$. Suppose that there exists $\bar{k}>0$ such that
$\mathcal{O}_{\bar{k}}=\{{\mathbf{x}}\in\mathbb{R}^{n}:0<V({\mathbf{x}})<\bar{k}\}\subset\mathcal{F}.$
Let ${\mathbf{x}}_{0}\in\mathcal{O}_{\bar{k}}$ with $V({\mathbf{x}}_{0})=k\leq\bar{k}$ and consider the maximal solution of $\dot{{\mathbf{x}}}=f({\mathbf{x}})+g({\mathbf{x}})u^{*}({\mathbf{x}})$ in $\mathcal{F}$ with initial condition at ${\mathbf{x}}_{0}$. We let ${\mathbf{x}}(\cdot;{\mathbf{x}}_{0})$ be such solution and $[0,T)$ its interval of existence (with either $T=\infty$ or $T<\infty$). Then, we have
$\lim\limits_{t\to T}{\mathbf{x}}(t;{\mathbf{x}}_{0})=\mathbf{0}_{n}.$
Proof.
Suppose first that there exists a compact subset $C\subset\mathcal{F}$ such that $x(t)\in C$ for all $t\in[0,T)$. Then, by [1, Proposition C.3.6], it follows that $T=\infty$. Since ${\mathbf{x}}(\cdot;{\mathbf{x}}_{0})$ is precompact, the omega limit set of ${\mathbf{x}}_{0}$, denoted as $\omega^{+}({\mathbf{x}}_{0})$ is non-empty. Since $V$ is a LaSalle function in $\mathcal{O}_{\bar{k}}$, for all $\boldsymbol{\xi}\in\omega^{+}({\mathbf{x}}_{0})$ we have $\dot{V}(\boldsymbol{\xi})=0$. However, since $\xi\in\omega^{+}({\mathbf{x}}_{0})\in\mathcal{O}_{\bar{k}}$, we have $\dot{V}(\boldsymbol{\xi})<0$, reaching a contradiction. Therefore, there exists no such set $C$. Let $\epsilon>0$ and consider the set $C_{\epsilon,k}=\{{\mathbf{x}}\in\mathbb{R}^{n}:\epsilon\leq V({\mathbf{x}})\leq k\}$. Since $C_{\epsilon,k}$ is a compact set contained in $\mathcal{F}$, there exists $t_{1}\in(0,T)$ such that ${\mathbf{x}}(t_{1};{\mathbf{x}}_{0})\notin C_{\epsilon,k}$. Let $t_{0}=\max_{s}\{s\leq t_{1}:{\mathbf{x}}(s;{\mathbf{x}}_{0})\in C_{\epsilon,k}\ \forall s\leq t\}$. First note that $V({\mathbf{x}}(t_{0};{\mathbf{x}}_{0}))=k$ is impossible. Indeed, for such a $t_{0}$ we have $\dot{V}({\mathbf{x}}(t_{0};{\mathbf{x}}_{0}))<0$, so $V({\mathbf{x}}(t;{\mathbf{x}}_{0}))>V({\mathbf{x}}(t_{1};{\mathbf{x}}_{0}))$ for all $t\in[t_{0}-\bar{\epsilon},t_{0}]$ for some small enough $\bar{\epsilon}>0$ (by continuity of $\dot{V}$), and therefore $V({\mathbf{x}}(t;{\mathbf{x}}_{0}))>k$ for $t\in[t_{0}-\bar{\epsilon},t_{0}]$, which contradicts the definition of $t_{0}$. Since $V({\mathbf{x}}(t_{0};{\mathbf{x}}_{0}))=k$ is impossible, we necessarily have $V({\mathbf{x}}(t_{0};{\mathbf{x}}_{0}))=\epsilon$. Since $\dot{V}({\mathbf{x}}(t;{\mathbf{x}}_{0}))<0$ for all $t\in[0,T)$, we conclude that $V({\mathbf{x}}(t;{\mathbf{x}}_{0}))<\epsilon$ for all $t\in(t_{0},T)$. Since $\epsilon$ is arbitrary, we necessarily have $\lim\limits_{t\to T}{\mathbf{x}}(t;{\mathbf{x}}_{0})=\mathbf{0}_{n}$. ∎
@@ 4 Neural Network Approximation of the Universal Formula
$\tilde{J}_{{\mathbf{q}}({\mathbf{p}})}({\mathbf{k}})$ $=-M\sum_{i=1}^{N}\frac{\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}}{M^{2}}+\frac{\lVert{\mathbf{k}}\rVert^{2}}{M^{2}}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}$
$=-\frac{1}{M}\sum_{i=1}^{N}\frac{\lVert{\mathbf{B}}_{i}\rVert^{2}+\lVert{\mathbf{k}}\rVert^{2}}{2(A_{i}+{\mathbf{B}}_{i}^{\top}{\mathbf{k}})}=J_{{\mathbf{p}}}({\mathbf{k}}).$
@@
Lemma 4.1 shows that by approximating $\tilde{k}^{*}$ with values of $\mathbf{q}$} in $\mathcal{T}=\left([-1,1]^{N}\times\mathcal{B}_{1}(\mathbf{0}_{m})^{N}\right)\cap\mathcal{P}\times[0,1]$, we can recover any value of $k^{*}$. Therefore, our approach consists in approximating the minimizer of $\tilde{J}_{{\mathbf{q}}}$ with a NN for $\mathbf{q}$} in the set $\mathcal{T}$. Since $\mathcal{P}$ is an open set, $\mathcal{T}$ is not compact. Given that the universal approximation theorem of neural networks [33] is only valid in compact sets, it does not apply to $\mathcal{T}$. However, for any compact subset of $\mathcal{T}$ (which we can take to be arbitrarily close to $\mathcal{T}$), the universal approximation theorem ensures that a NN with a sufficiently large number of parameters can approximate $k^{*}$ arbitrarily well. Note also that by choosing the activation functions of the NN to be smooth, the resulting approximation of $k^{*}$ is $\mathcal{C}^{\infty}(\mathcal{P})$, and therefore the approximation of the controller $u^{*}$ is $\mathcal{C}^{l}(\mathcal{P})$.
@@
(Applicability for multiple controllers): Note that once we have obtained an approximation of $k^{*}$, we can use it in many different instantiations of the functions $a_{1},\ldots,a_{N},b_{1},\ldots,b_{N}$, i.e., in a variety of different control problems. The NN only depends on the number of constraints NN and the dimension of the input $m$, but remarkably, does not depend on $n$, the dimension of the state. In fact, given a NN trained with a given input dimension $m$ and number of constraints $N$, all control problems with input dimension at most $m$ and at most $N$ constraints can be solved with the same NN (in case the input dimension is strictly less than $m$ or the number of constraints is strictly less than $N$ one can simply assign the coefficients of the additional inputs or constraints so that they are trivially satisfied). In particular, this implies the remarkable fact that all safe stabilization problems for systems with a given input dimension can be solved with the same NN approximation.  $\bullet$
@@
(Robustness): If one of the inequalities corresponds to a CLF inequality (2), and the corresponding CLF is an input-to-state stable (ISS) Lyapunov function for the system (1) (cf. [34, Section 3.3]), then if we let $\hat{k}$ be an approximated version of the controller $k^{*}$ (like the one obtained by approximating it with a NN), and the bound $\big\lVert\hat{k}({\mathbf{x}})-k^{*}({\mathbf{x}})\big\rVert\leq\sigma$ holds for all ${\mathbf{x}}\in\mathcal{X}$, with $\mathcal{X}$ some compact set, we have
$\nabla V({\mathbf{x}})^{\top}(f({\mathbf{x}})+g({\mathbf{x}})\hat{k}({\mathbf{x}}))$
$=\nabla V({\mathbf{x}})^{\top}(f({\mathbf{x}})+g({\mathbf{x}})k^{*}({\mathbf{x}}))+\gamma(\sigma),$
for some class $\mathcal{K}_{\infty}$ function $\gamma$. A similar result holds if the inequality is a CBF inequality (3) and the CBF defining it is an input-to-state safe (ISSf) barrier function for the system (1[35, Theorem 1]. Therefore, CLF and CBF inequalities are robust to approximation errors induced by the NN and therefore the NN-based controller, when applied to safety and stability tasks, makes the closed-loop system close to safe and stable, respectively.  $\bullet$
@@
(Hard-constrained NN): We should also point out that the recently introduced HardNet method [36] shows that, by including a projection step at the output layer of the NN, the NN predictions can be guaranteed to satisfy a set of affine constraints. Although the universal approximation guarantees are retained, the projection operation in the last layer makes the controller obtained from it locally Lipschitz but not $\mathcal{C}^{l}(\mathcal{F})$, for $l\geq 1$.  $\bullet$